OpenStudy (anonymous):

For the investment with interest compounded annually, find the final balance and the total interest earned. $500 for 2 years at 8% A. final balance: $540 interest earned: $40 B. final balance: $580 interest earned: $80 C. final balance: $583.20 interest earned: $83.20 D. final balance: $620 interest earned: $120

OpenStudy (jdoe0001):

\(\bf A=P\left(1+\frac{r}{n}\right)^{nt} \\ \quad \\ A=\textit{current amount}\\ P=\textit{original amount deposited}\to \$500\\ r=rate\to 8\%\to \frac{8}{100}\to 0.08\\ n=\textit{times it compounds per year, annually means once}\to 1\\ t=years\to 2 \\ \quad \\ A=500\left(1+\frac{0.08}{1}\right)^{1\cdot 2}\qquad \textit{final balance}\to P+A\)

OpenStudy (anonymous):

thanks

OpenStudy (jdoe0001):

yw